OpenStudy (anonymous):

(integral exponents) simplify; (9/16)^-2

OpenStudy (anonymous):

(9/16)^-2=(16/9)2 (9/16)^-2=256/81 (9/16)^-2=32 right or will it just be 256/81?

OpenStudy (gorv):

9=3^2 16=4^2 (3^2/4^2)^-2

OpenStudy (midhun.madhu1987):